OpenStudy (korosh23):

Physics 12 Question!

OpenStudy (korosh23):

The answer is right there, just please explain why is it tan 28, and not cos 28? Thank you.

OpenStudy (korosh23):

Also please draw me a free body diagram with the position of folcrum.

OpenStudy (anonymous):

I think what they did was divide the sum of the forces in y and x. So \[\huge \frac{ T_y }{ T_x }=\frac{ \sin(28) }{ \cos(28) }=\tan(28)\] The F_y and F_x that I put here are the total forces, they are only the components of that particular triangle (due to the cable) acting in those directions. Then, since they knew T_x they rearranged it so that\[\huge T_y=T_x \tan(28)\] \(T_y\) is the same as \(T_\perp\)

OpenStudy (anonymous):

`I think what they did was divide the sum of the forces in y and x` Sorry, NOT the sum of the forces, just the components of the tension!

OpenStudy (korosh23):

@CShrix I do not understand why did they do that for? can you explain more?